State Dalton’s law of partial pressure. Derive its mathematical form.
1
Expert's answer
2020-03-11T09:06:39-0400

Solution.

Partial is the pressure of a gas in a mixture that this gas would create in the absence of other components.

"p = p1 + p2 + p3 + ... + pn"

"p \times V = \nu \times R \times T"

"pn = \frac{\nu \times R \times T}{V}"

"p = \frac{R \times T}{V} \times (\nu1 + \nu2 + \nu3)"

Answer:

Partial is the pressure of a gas in a mixture that this gas would create in the absence of other components.

"p = \frac{R \times T}{V} \times (\nu1 + \nu2 + \nu3)"